# DirectorLLM for Human-Centric Video Generation

### v1

@@ Abstract
In this paper, we introduce DirectorLLM ***All data processing and model development for this research were conducted at Rutgers University, adhering to institutional guidelines., a novel video generation model that employs a large language model (LLM) as the “director” to orchestrate human poses within videos. As foundational text-to-video models rapidly evolve, the demand for high-quality human motion and interaction grows. To address this need and enhance the authenticity of human motions, we extend the LLM from a text generator to a video director and human motion simulator. Utilizing open-source resources from Llama 3, we train the DirectorLLM to generate detailed instructional signals, such as human poses, to guide video generation. This approach offloads the simulation of human motion from the video generator to the LLM, effectively creating informative outlines for human-centric scenes. These signals are used as conditions by the video renderer, facilitating more realistic and prompt-following video generation. As an independent LLM module, it can be applied to different video renderers, including UNet and DiT, with minimal effort. Experiments on automatic evaluation benchmarks and human evaluations show that our model outperforms existing ones in generating videos with higher human motion fidelity, improved prompt faithfulness, and enhanced rendered subject naturalness.
@@ 1 Introduction
The fields of image and video generation have witnessed remarkable advancements, largely due to the emergence of diffusion-based models and expansive text-to-image (T2I) frameworks like DALL-E 2 [31], Imagen [34], Stable Diffusion [33], Emu [9], eDiff-I [6], etc. These models empower users to create diverse and realistic visuals from text prompts. Image-conditioned models (Kandinsky [32], Stable Unclip [2]) provides enhanced fidelity to input images in generating variations. Building on these advances, video generation models such as VideoCrafter [8] and LaVie [42] have extended these frameworks to produce temporally consistent videos by incorporating temporal layers within the SD architecture. DiT [27] models, such as Sora [26], CogVideoX [17], and Movie Gen [29] have demonstrated unprecedented capacities, achieving longer and high-quality videos across various resolutions and aspect ratios. Methods like VideoComposer [41] and DreamVideo [44] integrate video generation with various control signals. Although these techniques significantly improve content generation and customization, they often face limitations in understanding human poses and motion dynamics across extended video durations.
@@
Recent advancements in text-to-image and text-to-video generation leverage large language models (LLMs) to improve the spatial and temporal coherence of generated content. Generating complex visual dynamics directly from text prompts remains a significant challenge for diffusion models, especially when maintaining coherent spatial and temporal relationships. To address this, recent research LVD  [21], LayoutGPT [11], attention-refocusing [28], and VideoDirectorGPT [22] have explored using LLMs like GPT-4 [1] and PaLM2 [4] to produce explicit spatiotemporal layouts, such as bounding boxes or frame-by-frame descriptions, derived from a single user prompt. These layouts then serve as a foundation for conditioning diffusion models, enhancing their ability to produce images and videos with coherent object relationships across frames. By transforming the narrative input into precise bounding boxes and layout structures, these methods rely on off-the-shelf LLM to simulate subject dynamics, improving motion consistency and dynamic realism.
@@
In this work, we present DirectorLLM, a novel approach to video generation from text prompts, designed to tackle the challenge of creating realistic and dynamic human-centric videos with improved human motions, especially those involving human interactions and intricate poses. Existing methods [8, 42] generally rely on diffusion models to synthesize videos directly from text prompts, requiring the model to simultaneously manage scene understanding, object interactions, and video rendering, often resulting in limited performance on complex scenarios. For human-centric videos, this causes anatomically inaccurate outputs, such as broken or extra limbs, unnatural joint movements, or abrupt shifts in human motion. To address these limitations, we introduce a multi-component model that decouples scene understanding and pose dynamics from the video generation process. Our approach consists of three main components: a specialized large language model (LLM) that interprets the text prompt to generate human poses at 1 frame per second (FPS), a compact linear diffusion model that interpolates these sparse poses to a smooth 30 FPS, enhancing temporal consistency, and a video generator (VideoCrafter [8]) enhanced with pose ControlNet [49] to render realistic videos conditioned on both the predicted human poses and text prompt, enabling it to render highly realistic and contextually coherent videos. This structured framework allows the LLM to handle scene understanding and human motion simulation, offloading these tasks from the video generator, which can then focus on creating visually accurate frames. By integrating these specialized components, our model achieves improved realism and coherence in human-centric video generation compared to prior methods. Inspired by real-world movie production, the overview of the proposed idea is depicted in Fig. 1 where director and cinematographer work together.
@@
Our contributions are as follows: (i) To our knowledge, this is the first work to train a large language model (e.g., Llama 3 [10]), VQ-VAE [38], and linear diffusion for simulating and generating human poses, specifically tailored for text-to-video generation task. (ii) We enhance diffusion-based text-to-video (T2V) synthesis by integrating predicted human motion, leading to improved realism and coherence in human-centric videos. (iii) Our system advances the current state of T2V synthesis, particularly in terms of human motion realism and temporal consistency.
2 Related Work
@@ 2.1 Text-to-Video (T2V) Generation
Text-to-video (T2V) generation leverages deep learning models to interpret text input and generate corresponding video content, building on earlier advancements in text-to-image generation by incorporating the added complexity of motion and temporal consistency. In recent years, T2V models have evolved rapidly, addressing challenges in capturing both spatial coherence and dynamic realism. Notable models, such as ModelScopeT2V [39] and LaVie [42], enhance video synthesis by integrating temporal layers within spatial frameworks to achieve more cohesive video output. VideoCrafter2 [8], for instance, tackles the scarcity of labeled video data by making use of high-quality image datasets to improve training. Transformer-based models like Sora [26] have achieved impressive results in generating videos with high visual quality, stable temporal consistency, and varied motion dynamics. Moreover, new advancements like Movie Gen [29] push the boundaries further, achieving exceptional quality in both video and audio generation, underscoring the rapid progress and potential of T2V models. Unlike existing text-to-video generators that simultaneously handle scene understanding and rendering, our model focuses on human-centric video generation, offloading high-level scene understanding and human pose simulation to a large language model, allowing the video generator to focus solely on rendering high-quality temporally consistent videos.
2.2 Motion-Aware T2V Models
Motion cues play a crucial role in video generation, enabling precise temporal motion control and enhancing the realism of synthesized videos. Recent works have explored incorporating explicit motion control by focusing on both camera and object movements. For camera motion control, methods like AnimateDiff [13], VideoComposer [41], CameraCtrl [14], Direct-A-Video [46], and MotionCtrl [43] design specific modules to encode camera trajectories, often relying on training with large-scale datasets, which leads to high computational costs. For object motion control, trajectory-based motion guidance is employed in works such as DragNUWA [47], MotionCtrl [43], Motion-I2V [35], Tora [51] and DragAnything [45], while box-based motion guidance is utilized by TrailBlazer [24] and Boximator [40]. These approaches typically perform control by training extra motion controllers on pre-trained video diffusion models. Additionally, methods like MotionDirector [52] use reference videos to extract motion patterns for generation but are limited to reproducing existing motions. These prior works leverage explicit motion signals, offering greater adaptability and user accessibility. In contrast to these models that require user-provided motion signals, our model generates human pose dynamics directly from text prompts using a fine-tuned large language model, enabling fully text-driven human-centric video generation.
2.3 LLM in Image and Video Generations
@@
Several recent text-to-image models have incorporated large language models (LLMs) to assist image and video generation. Methods like LayoutGPT [11] employs a program-guided method for layout-oriented visual planning across various domains. Additionally, Attention Refocusing [28] introduces innovative loss functions to realign attention maps according to specified layouts. Building upon these developments, recent efforts have extended the use of LLMs to assist text-to-video models. Free-bloom [18] leverages LLMs to generate detailed frame-by-frame descriptions from a single prompt, enriching the video’s narrative. LLM-grounded video generation [21] and VideoDirectorGPT [22] proposes a training-free approach to guide diffusion models using LLM-generated frame-by-frame layouts. GPT-4-Motion [23] uses GPT-4 [1] to generate scripts for Blender to produce scene components, which serve as conditions for Stable Diffusion to synthesize videos. Unlike previous work that use pre-trained and frozen language models to generate layout and bounding boxes from text prompts, our model fine-tunes a large language model, i.e., Llama 3 [10], to predict vector-quantized tokens to simulate human poses, enabling more accurate motion understanding tailored for video generation.
@@ 3.1 Preliminaries
Text-to-video. T2V diffusion models are designed to generate video content by adding noise in a latent space and then learning to denoise it. The process starts by encoding an input video into a latent representation, $\mathbf{z}_{0}$. Noise, represented by $\epsilon$, is incrementally added to this latent representation over time steps $t$, creating a noisy latent $\mathbf{z}_{t}$ that corresponds to different noise levels. This progression simulates a reverse Markov chain [33] process, where the diffusion model, $\epsilon_{\theta}$, is trained to predict and remove the noise added to $\mathbf{z}_{t}$ as it transitions back toward the clean latent representation. The model is trained using a reconstruction loss that minimizes the difference between the actual noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c},t)$, where $\mathbf{c}$ represents a conditional input, such as text or an image used for guidance. This is expressed mathematically as:
@@
Motion generation. Prior works on diffusion have explored audio to dance [37], text to motion [50, 19], audio to gestures [3, 5, 48] and speech to human pose [25]. Audio2Photoreal [25] presents a avatar pose generation systems conditioned on audio inputs. Given speech audio, it outputs multiple possibilities of gestural motion for an individual, including face, body, and hands. Unlike these pose models, our method focuses on generating dynamic human poses directly from text prompts at the context of text to video generation, leveraging a fine-tuned large language model for detailed and contextually accurate motion planning.
@@ 3.2 Methodology
Our model consists of three main stages: a large language model (e.g., tuned from Llama 3 [10]) generates discretized human pose tokens from text prompts at low FPS for human motion planning, followed by a linear diffusion model for pose interpolation. The final stage uses a video generation model (VideoCrafter2 [8]) enhanced by ControlNet [49] to produce realistic human-centric videos.
3.2.1 DirectorLLM
Our DirectorLLM, built on Llama 3 [10], plays a crucial role in scene understanding and human pose generation planning, serving as a high-level director that interprets the text prompt to produce structured, instance-level layouts for human motion. Given a prompt, the DirectorLLM identifies content of the requested video in the scene and predicts corresponding human pose tokens at 1 fps. To make these outputs compatible with our generative pipeline, we convert continuous human pose vectors into discrete tokens via a residual vector quantized autoencoder (VQ-VAE [38]), enabling the LLM to process human pose information as token sequences. Input sequences are $z=(t,p)$, in which $t=(t_{1},t_{2},\dots,t_{n})$ are text prompt tokens, and $p=(p_{1},p_{2},\dots,p_{n})$ are pose tokens. The model is trained on cross-entropy loss using the standard next-token prediction framework,
@@ 3.2.1 DirectorLLM
Figure 4: Examples of generated poses and videos frames from DirectorLLM. Poses are generated from the LLM and interpolated by pose diffusion model.
3.2.2 Diffusion Interpolator
We train a compact linear diffusion model to densify sparse poses generated by LLM into smooth and temporally consistent outputs at 30 FPS. As shown in Fig. 9 right part, this model enhances the initial sparse human poses by generating intermediate frames that maintain natural motion and dynamics. We incorporate conditional pose information through cross-attention layers, allowing the model to refine pose sequences frame-by-frame based on both structural and temporal cues. To improve fidelity further, we condition the model on CLIP [30] embeddings of the text prompt, enhancing alignment between the generated dynamics and the scene description. Additionally, we employ classifier-free guidance [16] by intermittently replacing conditions with zero tensors during training, promoting the model’s ability to synthesize realistic sequences even under partial guidance. This linear diffusion process enables seamless transition from sparse layouts to dense, high-frequency poses, preparing the input for the final video generation phase.
3.2.3 ControlNet Video Diffusion
In our video generator, we build our renderer on VideoCrafter2 [8], a text-to-video generator with a UNet architecture. We enhance following the design of ControlNet [49] on additional human pose signals, ensuring that it adheres closely to the predicted human poses and dynamics throughout the video generation process. We incorporate zero convolution layers between ControlNet and original UNet, which are initialized with zero weights, minimizing its interference with the original U-Net’s operations at the beginning of training.
We choose a UNet structured model as our base video generator instead of a diffusion transformer (DiT) due to the substantial computational demands of DiT. Diffusion transformer requires self-attention across all spatial and temporal locations, which makes it very costly. In our design, we offload the high-level motion understanding and planning a dedicated large language model (LLM) and a pose interpolator, reducing the workload for the video generator. This division of tasks enables the UNet model to focus exclusively on rendering details based on the human poses generated by the prior components, appropriate for generating realistic, temporally consistent human-centric video frames, making it a efficient and suitable choice.
@@ 4.1 Dataset
For data preparation, we start data preparation with the Shutterstock (SSTK) [36] video dataset, generating detailed text prompts with a video captioning model to provide scene descriptions. As our task focus on human-centric video generation, we filter the dataset according to text prompt and only keep human-centric video samples. We collected videos where the same person appears across all frames following [29]. Then we apply OpenPose [7] on each frame to extract precise pose information. The extracted poses have 18 key points, disregrading all hand and feet points. We limits a maximum sequence length of 240 frames to capture extended scenes. During training, a subsequence of 200 frames is randomly sampled, with shorter sequences zero-padded to reach this length. From the 200 complete frames, we evenly extract 20 key frames. This thorough filtering and annotation process produces a high-quality dataset, well-suited for our model’s structured scene planning and pose conditioning tasks.
The prompts generated by our video caption model are highly structured, with the first sentence providing an overall video description and the following sentences offering detailed descriptions of appearance and scene elements. We call the first sentence “Subject-Level Prompt” as shown in Fig. 1, as it contains the key information about human motions and poses. As detailed descriptions are less related to human pose, we only use Subject-Level Prompt as input in our LLM and pose interpolator.

### v2

# Llama Learns to Direct: DirectorLLM for Human-Centric Video Generation

@@ Abstract
In this paper, we introduce DirectorLLM, a novel video generation model that employs a large language model (LLM) as the “director” to simulate human poses within videos. To enhance the authenticity of human motions in text-to-video models, we extend the LLM from a text generator to a video director and human motion simulator. We train the DirectorLLM to generate detailed human poses, to guide video generation, offloading the simulation of human motion from the video generator to an LLM, effectively creating informative outlines for human-centric scenes. These signals are used as conditions by the video renderer, facilitating more realistic and prompt-following video generation. Experiments on automatic evaluation benchmarks and human evaluations show that our model outperforms existing ones in generating videos with higher human motion fidelity, improved prompt faithfulness, and enhanced rendered subject naturalness.
1 Introduction
@@
The fields of image and video generation have witnessed remarkable advancements, largely due to the emergence of diffusion-based models like DALL-E 2 Ramesh et al. (2022), Imagen Saharia et al. (2022), Stable Diffusion Rombach et al. (2022), Emu Dai et al. (2023), eDiff-I Balaji et al. (2022), Kandinsky Razzhigaev et al. (2023), Stable Unclip AI (2024). These models empower users to create diverse and realistic visuals from text prompts. Building on these advances, video generation models such as VideoCrafter Chen et al. (2023) and LaVie Wang et al. (2023b), Sora OpenAI (2024), CogVideoX Hong et al. (2022), and Movie Gen Polyak et al. (2024) have demonstrated unprecedented capacities, achieving longer and high-quality videos across various resolutions and aspect ratios. Methods like VideoComposer Wang et al. (2024b) and DreamVideo Wei et al. (2024) integrate video generation with various control signals. Although these techniques significantly improve content generation and customization, they often face limitations in understanding human poses and motion dynamics across extended video durations.
@@
Recent advancements in text-to-image and text-to-video generation leverage large language models (LLMs) to improve the spatial and temporal coherence of generated content. Generating complex visual dynamics directly from text prompts remains a significant challenge for diffusion models, especially when maintaining coherent spatial and temporal relationships. To address this, recent research LVD Lian et al. (2023), LayoutGPT Feng et al. (2024), attention-refocusing Phung et al. (2024), and VideoDirectorGPT Lin et al. (2023) have explored using LLMs like GPT-4 Achiam et al. (2023) and PaLM2 Anil et al. (2023) to produce explicit spatiotemporal layouts, such as bounding boxes or frame-by-frame descriptions, derived from a single user prompt. These layouts then serve as a foundation for conditioning diffusion models, enhancing their ability to produce images and videos with coherent object relationships across frames. By transforming the narrative input into precise bounding boxes and layout structures, these methods rely on off-the-shelf LLM to simulate subject dynamics, improving motion consistency and dynamic realism.
@@
In this work, we present DirectorLLM, a novel T2V model designed to tackle the challenge of creating realistic and dynamic human-centric videos with improved human motions. Existing methods Chen et al. (2023); Wang et al. (2023b) generally rely on diffusion models to synthesize videos directly from text prompts, requiring the model to simultaneously manage scene understanding, object interactions, and video rendering, often resulting in anatomically inaccurate human body outputs, such as broken or extra limbs, unnatural joint movements, or abrupt shifts in human motion. To address these limitations, we decouple scene understanding and pose simulation from the video generation process. Our approach consists of three main components: a specialized large language model (LLM) that interprets the text prompt to generate human poses at 1 frame per second (FPS), a compact linear diffusion model that interpolates these sparse poses to a smooth 30 FPS, enhancing temporal consistency, and a video generator (VideoCrafter Chen et al. (2023)) enhanced with pose ControlNet Zhang et al. (2023) to render realistic videos conditioned on both the predicted human poses and text prompt . This structured framework allows the LLM to handle scene understanding and human motion simulation, offloading these tasks from the video generator, which can then focus on creating visually accurate frames. By integrating these specialized components, our model achieves improved realism and coherence in human-centric video generation compared to prior methods.
@@
Our contributions are as follows: (i) To our knowledge, this is the first work to train a large language model (e.g., Llama 3 Dubey et al. (2024)), residual VQ-VAE Van Den Oord et al. (2017), and linear diffusion for simulating and generating human poses, specifically tailored for text-to-video generation task. (ii) We enhance diffusion-based text-to-video (T2V) synthesis by integrating predicted human motion, leading to improved realism and coherence in human-centric videos. (iii) Our system advances the current state of T2V synthesis, particularly in terms of human motion realism and temporal consistency.
@@ 1.1 Related Work
Text-to-video (T2V) generation. Text-to-video (T2V) generation leverages deep learning models to interpret text input and generate video content, and it’s been evolving rapidly. Notable models, such as ModelScopeT2V Wang et al. (2023a) and LaVie Wang et al. (2023b), enhance video synthesis by integrating temporal layers within spatial frameworks to achieve more cohesive video output. VideoCrafter2 Chen et al. (2023), for instance, tackles the scarcity of labeled video data by making use of high-quality image datasets to improve training. Transformer-based models like Sora OpenAI (2024) have achieved impressive results in generating videos with high visual quality, stable temporal consistency, and varied motion dynamics. Moreover, Movie Gen Polyak et al. (2024) pushes the boundaries further, achieving exceptional quality in both video and audio generation. Recent work has explored the incorporation of explicit motion controls. Notable works include AnimateDiff Guo et al. (2023), VideoComposer Wang et al. (2024b), CameraCtrl He et al. (2024a), Direct-A-Video Yang et al. (2024), and MotionCtrl Wang et al. (2024c), which design specific modules to encode camera trajectories. Trajectory-based motion guidance is used in works such as DragNUWA Yin et al. (2023), MotionCtrl Wang et al. (2024c), Motion-I2V Shi et al. (2024), Tora Zhang et al. (2024), and DragAnything Wu et al. (2025), while box-based motion guidance is used by TrailBlazer Ma et al. (2023) and Boximator Wang et al. (2024a). Additionally, methods like MotionDirector Zhao et al. (2023) use reference videos to extract motion patterns for generation. In contrast to these models, which require user-provided motion signals, our model takes care of generating human poses. , enabling fully text-driven human-centric video generation.
@@
LLM in image and video generation. Several recent text-to-image models have incorporated large language models (LLMs) to assist image and video generation. Methods like LayoutGPT Feng et al. (2024) employ a program-guided method for layout-oriented visual planning across various domains. Additionally, Attention Refocusing Phung et al. (2024) introduces innovative loss functions to realign attention maps according to specified layouts. Building upon these developments, recent efforts have extended the use of LLMs to assist text-to-video models. Free-bloom Huang et al. (2024) leverages LLMs to generate detailed frame-by-frame descriptions from a single prompt, enriching the video’s narrative. LLM-grounded video generation Lian et al. (2023) and VideoDirectorGPT Lin et al. (2023) propose a training-free approach to guide diffusion models using LLM-generated frame-by-frame layouts. GPT-4-Motion Lv et al. (2024) uses GPT-4 Achiam et al. (2023) to generate scripts for Blender to produce scene components, which serve as conditions for Stable Diffusion to synthesize videos. Unlike previous work that uses pre-trained and frozen language models to generate layout and bounding boxes from text prompts, our model fine-tunes a large language model, Llama 3 Dubey et al. (2024), to predict vector-quantized tokens to simulate human poses, enabling more accurate motion understanding tailored for video generation.
@@ 2.1 Preliminaries
Text-to-video. T2V diffusion models are designed to generate video content by adding noise in a latent space and then learning to denoise it. The process starts by encoding an input video into a latent representation, $\mathbf{z}_{0}$. Noise, represented by $\epsilon$, is incrementally added to this latent representation over time steps $t$, creating a noisy latent $\mathbf{z}_{t}$ that corresponds to different noise levels. This progression simulates a reverse Markov chain Rombach et al. (2022) process, where the diffusion model, $\epsilon_{\theta}$, is trained to predict and remove the noise added to $\mathbf{z}_{t}$ as it transitions back toward the clean latent representation. The model is trained using a reconstruction loss that minimizes the difference between the actual noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c},t)$, where $\mathbf{c}$ represents a conditional input, such as text or an image used for guidance. This is expressed mathematically as:
@@
Motion generation. Prior works on diffusion have explored audio to dance Tseng et al. (2023), text to motion Zhang et al. (2022); Jiang et al. (2023), audio to gestures Alexanderson et al. (2023); Ao et al. (2023); Yu et al. (2023) and speech to human pose Ng et al. (2024). Audio2Photoreal Ng et al. (2024) presents a avatar pose generation systems conditioned on audio inputs. Given speech audio, it outputs multiple possibilities of gestural motion for an individual, including face, body, and hands. Unlike these pose models, our method focuses on generating dynamic human poses directly from text prompts at the context of text to video generation, leveraging a fine-tuned large language model for detailed and contextually accurate motion planning.
@@ 2.2 Methodology
Our model consists of three main stages: a large language model (e.g., tuned from Llama 3 Dubey et al. (2024)) generates discretized human pose tokens from text prompts at low FPS for human motion planning, followed by a linear diffusion model for pose interpolation. The final stage uses a video generation model (VideoCrafter2 Chen et al. (2023)) enhanced by ControlNet Zhang et al. (2023) to produce realistic human-centric videos.
DirectorLLM. Our DirectorLLM, built on Llama 3 Dubey et al. (2024), plays a crucial role in scene understanding and human pose generation planning, serving as a high-level director that interprets the text prompt to produce structured, instance-level layouts for human motion. Given a prompt, the DirectorLLM identifies content of the requested video in the scene and predicts corresponding human pose tokens at 1 fps. To make these outputs compatible with our generative pipeline, we convert continuous human pose vectors into discrete tokens via a residual vector quantized autoencoder (VQ-VAE Van Den Oord et al. (2017)), enabling the LLM to process human pose information as token sequences. Input sequences are $z=(t,p)$, in which $t=(t_{1},t_{2},\dots,t_{n})$ are text prompt tokens, and $p=(p_{1},p_{2},\dots,p_{n})$ are pose tokens. The model is trained on cross-entropy loss using the standard next-token prediction framework,
@@
Figure 4: Examples of generated poses and videos frames from DirectorLLM. Poses are generated from the LLM and interpolated by pose diffusion model. We show more pose-video sequences in the supplementary material.
Diffusion interpolator. We train a compact linear diffusion model to densify sparse poses generated by LLM into smooth and temporally consistent outputs at 30 FPS. As shown in Fig. 8 right part, this model enhances the initial sparse human poses by generating intermediate frames that maintain natural motion and dynamics. We incorporate conditional pose information through cross-attention layers, allowing the model to refine pose sequences frame-by-frame based on both structural and temporal cues. To improve fidelity further, we condition the model on CLIP Radford et al. (2021) embeddings of the text prompt, enhancing alignment between the generated dynamics and the scene description. Additionally, we employ classifier-free guidance Ho and Salimans (2022) by intermittently replacing conditions with zero tensors during training, promoting the model’s ability to synthesize realistic sequences even under partial guidance. This linear diffusion process enables seamless transition from sparse layouts to dense, high-frequency poses, preparing the input for the final video generation phase.
ControlNet video diffusion. In our video generator, we build our renderer on VideoCrafter2 Chen et al. (2023), a text-to-video generator with a UNet architecture. We enhance following the design of ControlNet Zhang et al. (2023) on additional human pose signals, ensuring that it adheres closely to the predicted human poses and dynamics throughout the video generation process. We incorporate zero convolution layers between ControlNet and original UNet, which are initialized with zero weights, minimizing its interference with the original U-Net’s operations at the beginning of training. We choose a UNet structured model as our base video generator instead of a diffusion transformer (DiT) due to the substantial computational demands of DiT. Diffusion transformer requires self-attention across all spatial and temporal locations, which makes it very costly. In our design, we offload the high-level motion understanding and planning a dedicated large language model (LLM) and a pose interpolator, reducing the workload for the video generator. This division of tasks enables the UNet model to focus exclusively on rendering details based on the human poses generated by the prior components, appropriate for generating realistic, temporally consistent human-centric video frames, making it a efficient and suitable choice.
@@ 3.1 Dataset
For data preparation, we start data preparation with the Shutterstock (SSTK) Shutterstock (2024) video dataset, generating detailed text prompts with a video captioning model to provide scene descriptions. As our task focus on human-centric video generation, we filter the dataset according to text prompt and only keep human-centric video samples. We collected videos where the same person appears across all frames following Polyak et al. (2024). Then we apply OpenPose Cao et al. (2017) on each frame to extract precise pose information. The extracted poses have 18 key points, disregrading all hand and feet points. We limits a maximum sequence length of 240 frames to capture extended scenes. During training, a subsequence of 200 frames is randomly sampled, with shorter sequences zero-padded to reach this length. From the 200 complete frames, we evenly extract 20 key frames. This thorough filtering and annotation process produces a high-quality dataset, well-suited for our model’s structured scene planning and pose conditioning tasks.